OpenStudy (anonymous):

Medal and fan!! Which ratios are proportional to 14 : 21? Choose all answers that are correct. A. 2 : 3 B. 3 : 2 C. 28 : 63 D. 56 : 84

OpenStudy (anonymous):

A cuz 14/7=2 and 21/7=3

OpenStudy (anonymous):

and

OpenStudy (anonymous):

2:3 is your answer as @roz_mar23 said

OpenStudy (anonymous):

but it says choose all that are correct

OpenStudy (unanimoose):

There won't always be multiple answers.

OpenStudy (anonymous):

14*4/21*4 = 56/84

OpenStudy (emmamink):

2:3

OpenStudy (anonymous):

so A and D

OpenStudy (anonymous):

yup